OpenStudy (anonymous):

Find the height of a cylinder with radius 2.5 cm and surface area 88.59 cm2.

Directrix (directrix):

The surface area of a cylinder is found by the following: Circumference of Base times the height plus 2 times the area of the base. SA = 2*pi*r*h + 2* pi*r^2 88.59 = 2*pi*2.5*h + 2*pi* (2.5)^2 It's all yours, @dopegyal_

OpenStudy (anonymous):

3.1

OpenStudy (anonymous):

?

Directrix (directrix):

The surface area of a cylinder is found by the following: Circumference of Base times the height plus 2 times the area of the base. SA = 2*pi*r*h + 2* pi*r^2 88.59 = 2*pi*(2.5)*h + 2* pi*(2.5)^2

Directrix (directrix):

Yes, I got what you did, 3.13 which is 3.1 to the nearest tenth.